OpenStudy (anonymous):

Find the values of the trigonometric functions of t from the given information. tan t= -(12/5), cos(t) >0 Find sin, cos, csc, sec, cot (T)

OpenStudy (campbell_st):

this is a 4h quadrant angle since tan is negative and cos is positive so draw a diagram  find h using pythagoras' theorem and then you'll be able to find all your trig ratios

OpenStudy (anonymous):

Given this, I got h=sqrt119 Then sin would have been (-12sqrt119)/119. However, this was given an incorrect.

OpenStudy (campbell_st):

well no, its h^2 = (5)^2 + (-12)^2 h^2 = 25 + 144 h^2 = 169 so h = 13 hope this helps

OpenStudy (anonymous):

Greatly. Thank you!